OpenStudy (anonymous):

MEDAL FOR BEST ANSWER Please walk through it with me Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 17 feet by 17 feet by 4 inches for a patio if the concrete costs $52.00 per cubic yard?

OpenStudy (anonymous):

If it is in cubic yard, we are dealing with a cube Volume of Cube = (Base)(Height)(Width)

OpenStudy (anonymous):

They give us all the dimensions for this Volume of Cube = (17)(17)(4) What is that?

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Did you get 1156?

OpenStudy (anonymous):

1156

OpenStudy (anonymous):

Okay perfect. So now you know that there are 1156 cubic yards. The second part of the question tells us that each cubic yard is $52 1 cubic yard = $52 1156 cubic yards = $? You have to multiply 52 by 1156 to get your answer.

OpenStudy (anonymous):

60112 ???

OpenStudy (anonymous):

Correct :) It costs that many dollars to pour the slab

OpenStudy (anonymous):

But on my study guide that has answers that isnt an answer choice.

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (anonymous):

What are the answer choices?

OpenStudy (anonymous):

The correct answer should be 185.53

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (anonymous):

Oh ha ha I see the problem. The dimensions were given in feet, not yards. You have to convert it.

OpenStudy (anonymous):

Ugh How do you do that?

OpenStudy (anonymous):

1 yard = 3 feet 17 feet = 5.666 yards

OpenStudy (anonymous):

soo 5.6 x 5.6x14?

OpenStudy (anonymous):

4 no 14

OpenStudy (anonymous):

The 4 is actually in inches 1 yard = 36 inches

OpenStudy (anonymous):

= 1/9 yards

OpenStudy (anonymous):

This isnt making sense..

OpenStudy (anonymous):

Multiply (5.6)(5.6)(1/9) = 3.567

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

Multiply that by 52 dollars, and you get 185.507

OpenStudy (anonymous):

It's off depending on how many 0.66666 you do

OpenStudy (anonymous):

i got 181

OpenStudy (anonymous):

Yeah you have to do (5.66666666666)(5.666666666666)(1/9) to be more accurate

OpenStudy (anonymous):

how many 6's is that

OpenStudy (anonymous):

Just type a lot. It is an infinite amount, but the more you do the closer to the right answer you will have.